The management of Stag Inc. was reviewing its equipment for impairment. The equipment had a cost of $900,000 with depreciation to date of $400,000 as of December 31, 2019. At this date the management has projected the present value of the future cash flows from the equipment to be $300,000. An equipment appraiser indicated that the equipment would likely sell for $322,000 net of a 15% transaction fee.
Stag Inc. intends to continue using the equipment in the future and estimates that it still has four years of useful life remaining. Straight line depreciation is applied. The financial year end December 31.
Required:
a) Perform the impairment test on the equipment.
b) Prepare the journal entry (if any) to record the impairment at December 31, 2019.

c) Calculate the annual depreciation for 2020.
Answer #1

A) Impairment test on the equipment:

This equipment will be recorded at Recoverable Amount.

Recoverable Amount will be the higher of

a)Fair value less cost to sell = 322000 - 15% of 322000 = 273700

b)Value in use = Present value of Future cash flows = 300000

So, Recoverable amount = 300000

Carrying amount of Asset = 900000-400000 = 500000

which is greater than Recoverable Amount,

Therefore, Impairment loss to be recognised.

B) impairment = Carrying Amount - Recoverable Amount = 500000-300000 = 200000

Journal Entry to record the impairment at December 31, 2019.

Account title Debit($) Credit($)
Impairment Loss 200000
To Equipment 200000

C) Annual depreciation for 2020:

Annual Depreciation = 300000/4 = 75000

Account title Debit($) Credit($)
Depreciation 75000
To Equipment 75000